OpenStudy (anonymous):

1.Simplify the following expression x - 4y + 2z - 3x - 3y + 2z = ?. 2.Evaluate the following expression using the values given: Find 4x - y - 2z if x = -2, y = 3, and z = -2

OpenStudy (texaschic101):

x - 4y + 2z - 3x - 3y + 2z --- lets put our like terms together x - 3x = -2x -4y - 3y = -7y 2z + 2z = 4z now put it all together and you get : -2x - 7y + 4z ====================== 4x - y - 2z ....when x = -2, y = 3 and z = -2 time to sub 4(-2) - 3 - 2(-2) -8 - 3 + 4 = - 7 =========== any questions ?

OpenStudy (anonymous):

when you said put it all together and you get -2x - 7y + 4z wouldn't it be 2z + 2z = 4z ?

OpenStudy (anonymous):

-2x + 7y + 4z -2x -7y + 4z -3x + 7y + 4z -3y - 7z + 4z